# Interpreting Equivariant Representations

### v1

@@ Appendix C Rotation invariant MNIST-classifier
The SO(2) invariant MNIST classifier is constructed using the tools provided in the ESCNN library provided by Weiler & Cesa (2019); Cesa et al. (2022b). That is, the model consists of 6 SO(2) steerable planar convolutions each of which is followed by batch-normalization and the FourierELU activation function. Each steerable planar convolution uses two irreducible representations to describe the output type; one invariant and one equivariant. The first two planar convolutions contains 16 feature maps, the next two 32 feature maps and the last two 64 feature maps. After each pair of planar convolutions a pooling layer is applied, and after the last convolution pooling is done over the spatial dimension to ensure invariance. Lastly, an invariant classifier is appended to the model. This is implemented using specifying a convolution using a kernel size of $1\times 1$, and using the trivial representation to describe the output type, followed by a fully-connected classification network.
@@
The model was trained using a cross-entropy loss. A learning rate of 0.01 and a batch-size of 128 was chosen. The model was trained for 100 epochs. The MNIST dataset was obtained through the pytorch libraryPaszke et al. (2017).

### v2

@@ Appendix C Rotation invariant MNIST-classifier
The SO(2) invariant MNIST classifier is constructed using the tools provided in the ESCNN library provided by (Weiler & Cesa, 2019; Cesa et al., 2022b). That is, the model consists of 6 SO(2) steerable planar convolutions each of which is followed by batch-normalization and the FourierELU activation function. Each steerable planar convolution uses two irreducible representations to describe the output type; one invariant and one equivariant. The first two planar convolutions contains 16 feature maps, the next two 32 feature maps and the last two 64 feature maps. After each pair of planar convolutions, a pooling layer is applied, and after the last convolution pooling is done over the spatial dimension to ensure invariance. Lastly, an invariant classifier is appended to the model. This is implemented using specifying a convolution using a kernel size of $1\times 1$, and using the trivial representation to describe the output type, followed by a fully-connected classification network.
@@
The model was trained using a cross-entropy loss. A learning rate of 0.01 and a batch-size of 128 was chosen. The model was trained for 100 epochs. The MNIST dataset was obtained through the pytorch library(Paszke et al., 2017).